Glycerol, C3H8O3, is a substance used extensively in the manufacture of cosmetics, foodstuffs, antifreeze, and plastics. Glycerol is a water-soluble liquid with a density of 1.2656 g/mL at 15°C. Calculate the molarity of a solution of glycerol made by dissolving 80.000 mL glycerol at 15°C in enough water to make 350.00 mL of solution. M
Answer #1

mass of C3H8O3 = density * volume

= 1.2656 g/mL * 80.000 mL

= 101.248 g

Molar mass of C3H8O3,

MM = 3*MM(C) + 8*MM(H) + 3*MM(O)

= 3*12.01 + 8*1.008 + 3*16.0

= 92.094 g/mol

mass(C3H8O3)= 101.248 g

use:

number of mol of C3H8O3,

n = mass of C3H8O3/molar mass of C3H8O3

=(1.012*10^2 g)/(92.09 g/mol)

= 1.099 mol

volume , V = 3.5*10^2 mL

= 0.35 L

use:

Molarity,

M = number of mol / volume in L

= 1.099/0.35

= 3.141 M

Answer: 3.14 M